y = 4 – 2x ?A.
x-intercept = 4; y-intercept = -2
B.
x-intercept = -2; y-intercept = 4
C.
x-intercept = 2; y-intercept = 4
D.
x-intercept = 4; y-intercept = 2

Answer by ewatrrr(24785)


Hi,
*Note: general slope-intercept form for the equation of a line:
y = mx + b where m is the slope and b the value of y where the line intercepts the y-xis
y = 4 - 2x Or y = -2x + 4 therefore, y-intercept is 4 (the value of y when x = 0)
x- intercept is when y = 0
0 = 4 - 2x
x = 2
answer C would be the correct choice